OpenStudy (lmooosey):

Use DeMoivre's Theorem to find the following. Write your answer in standard form. (sqrt2 cis 9pi/20)^5

OpenStudy (lmooosey):

@ProHeIper

OpenStudy (holsteremission):

\[\left(\sqrt2~\mathrm{cis}\,\frac{9\pi}{20}\right)^5=(\sqrt2)^5~\mathrm{cis}\,\frac{45\pi}{20}=\cdots \] since DeMoivre's theorem says that \[(\mathrm{cis}\,t)^n=\mathrm{cis}\,nt\]